OpenStudy (anonymous):

Which of the following is equivalent to x2 < -2x? x x x x -2 0 -2 -1

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Well, you can divide 'x' to both sides. \(\sf x^2 < -2x\) \(\sf \dfrac{x^2}{x} < \dfrac{-2}{x}\)

OpenStudy (igreen):

It cancels on the right, giving us: \(\sf \dfrac{x^2}{x} < -2\) Now simplify the left side

OpenStudy (anonymous):

please I still do not understand @iGreen

OpenStudy (igreen):

\(\sf x^2 \rightarrow x \cdot x\) So we have: \(\sf \dfrac{x \cdot x}{x} < -2\) Cancel out one of the \(\sf x\)'s on the top and one of the \(\sf x\)'s on the bottom.

OpenStudy (anonymous):

so is it x ?

OpenStudy (igreen):

Yes, it just leaves us with 'x'. So our final answer is: \(\sf x < -2\)

OpenStudy (anonymous):

@iGreen thank you so much

OpenStudy (igreen):

Np